Abcd is a parallelogram angle a is 2x plus 50 b is 3 x plus 40 find the measure of x find the measure of all 4
Answer by checkley77(12844)

2(2x+50)+2(3x+40)=360
4x+100+6x+80=360
10x=360-100-80
10x=180
x=180/10
x=18 degrees .
2*18+50
36+50=86 degrees for the smaller angle.
3*18+40
54+40=94 degrees for the larger angle.
Proof:
2*86+2*98=360
172+188=360
360=360